1. A U.S. corporation has purchased currency call options to hedge a £75,000 payable in 9 months. The premium is $.02 and the exercise price of the option is $1.50. If the spot rate at the time of maturity is $1.65, what is the total amount paid by the corporation if it acts rationally?

2.A U.S. corporation has purchased currency put options to hedge a £150,000 receivable in 9 months. The premium is $.02 and the exercise price of the option is $1.50. If the spot rate at the time of maturity is $1.65, what is the total amount received by the corporation if it acts rationally?
Answer #1

Payable amount= £75,000  
Excercise price of option =$   1.5      
Premium paid=    0.02      
Spot rate at maturity =$   1.65      


Value of call option = Spot rate - Excercise price          
Net profit or loss = Value of call option - Premium paid          


If Spot rate is less than Excercise price, call option will not be excercised. Value shall be nil. in this case loss will be equal to premium paid.           
VC = 1.65-1.50=   0.15      
Net profit =1.1-0.02=   0.13      
          
Total amount paid for payables = Payable amount *(Spot rate - Profit)          
75000*(1.65-0.13)          
114000          
          
So a rationale corporation will excercise option and Net amount payable is    $114,000.00      
          
          
          
Receivable amount= £150000  
Excercise price of put option =$   1.5      
Premium paid=   0.02      
Spot rate at maturity =$   1.65      


Value of put option = Excercise price-Spot rate          
Net profit or loss = Value of put option - Premium paid          


If Spot rate is more than Excercise price, put option will not be excercised. Value shall be nil. in this case loss will be equal to premium paid.           
VP = 1.50-1.65=   0  

Option will not be excercised  
Net profit =0-0.02=   -0.02      
          
Total amount received for receivables = receivables amount *(Spot rate +profit-loss)          
150000*(1.65-0.02)          
=244500          
          
So a rationale corporation will not excercise option and Net amount receivables is   $244,500.00      
          

(Please thumbs up or post comments if any query)